Operations on​ Functions:
2. Given the following​ functions, find each function combination. Write the new functions in descending order.
f(x)=x^2−3 and g(x)=x+5
a.(f+g)(x)=___________
b.(f−g)(x)=___________
c.(fg)(x)=__________
d.(f/g)(x)=________

Answer by josgarithmetic(39629)

------------------------------
f(x)=x^2−3 and g(x)=x+5
a.(f+g)(x)=___________
b.(f−g)(x)=___________
c.(fg)(x)=__________
d.(f/g)(x)=________
-----------------------------


(a)
Add the two functions.

(b)
Subtract g from f.

(c)
Multiply the two functions and simplify the product.

(d)
Try using polynomial division...